# Normalized Conditional Mutual Information Surrogate Loss for Deep Neural Classifiers

### v2

@@ I Introduction
Cross entropy (CE), first introduced by Cox [Cox_CE] for binary classification in the 1950s as the objective function for analyzing binary sequences, has become the de facto objective function for the most modern supervised learning algorithm for classification, while its capability for multi-class classification has not been justified until very recently [CMI]. In practice, CE minimizes the negative log-likelihood with one-hot targets, which geometrically pulls predictions toward the corners of the simplex.
@@
A large body of work augments CE with additional regularizers or proposes ad-hoc alternatives that empirically compete with CE. Specifically, orthogonal projection loss (OPL) [ranasinghe2021orthogonal] augments CE by maximizing cosine similarity among intra-class features while driving inter-class similarities toward zero; Squentropy [hui2023cut] applies an $\ell_{2}$ penalty to non–ground-truth entries of the output probability distribution, supervised contrastive learning (SupCon) [khosla2020supervised] extends the infoNCE [oord2018representation] loss to the supervised setting by treating all samples in the batch samples sharing the anchor’s label as positives, and SquareLoss [hui2021evaluation] argues that mean-squared-error (MSE) loss, with careful hyperparameter tuning, can match or even outperform CE. However, these approaches typically remain CE-centric; their performance depends on CE, while the auxiliary terms are ineffective on their own, require extensive hyperparameter tuning, or rely on very large batch sizes to realize their gains.
We note that in previous approaches, the learning process can be regarded as optimizing the model so that the output clusters are pulled toward predefined, corresponding distributions on the probability simplex. In this paper, we aim to answer the following question:
Instead of pulling output toward pre-defined target distributions, can we facilitate the learning by shaping the output distribution to be concentrated within classes and well-separated across classes using an information-theoretic principle?
To this end, following [CMI], we model the DNN-based classification task as a three-state Markov chain, as illustrated in Figure˜1. We quantify the concentration of the output distribution using the conditional mutual information (CMI) between the input and output, conditioned on the ground-truth label, and we quantify separation with $\Gamma$ (see Section˜III-B). We then define the normalized conditional mutual information (NCMI) as the ratio between the CMI and $\Gamma$. During training, instead of maximizing the log-likelihood (i.e., minimizing CE), we train the model to minimize its NCMI. After training, we evaluate the model using centroid-based decisions by comparing its outputs to class centroids, and via linear probing. Empirically, NCMI outperforms prior state-of-the-art losses and serves as an efficient surrogate for classification. The key contributions are as follows:
$\bullet$ We propose a new CE alternative named NCMI, for training DNN-based classifiers.
$\bullet$ To minimize the NCMI, we introduce a novel alternating optimization algorithm that minimizes the NCMI loss.
$\bullet$ To evaluate the NCMI’s effectiveness, we conduct comprehensive experiments on two natural-image datasets, namely, CIFAR-100 [krizhevsky2009learning] ImageNet [5206848] and two whole-slide image (WSI) datasets, CAMELYON-17 [bandi2018detection] and BRACS [brancati2022bracs]. Although modern DNNs and optimization tricks are tailored to CE–based surrogates, NCMI achieves state-of-the-art classification performance across all benchmarks.
@@ II Related work
Another line of work improves classification accuracy by modifying CE. Focal Loss [lin2017focal] down-weights well-classified examples via a power transformation so training emphasizes hard instances. PolyLoss [leng2022polyloss] reframes standard classification losses as polynomial expansions. Hui et al. propose SquareLoss [hui2021evaluation], and empirically found that squared loss performs on par with or even outperforms CE on modern DNNs. Supervised contrastive learning (SupCon) [khosla2020supervised] pulls together same-class embeddings and pushes apart different-class embeddings, followed by a linear classifier trained on the frozen features. Further, to mitigate overconfident predictions, label smoothing (LS) [szegedy2016rethinking] softens the one-hot targets, which can inadvertently produce compact class clusters [muller2019does]. AntiClass [kat2025anticlasses] replaces the one-hot target with a one-cold target to mitigate neural collapse.
In contrast, this paper studies the surrogate loss for classification through the lens of information geometry. We view a DNN as a mapping from $\boldsymbol{x}\in\mathbb{R}^{d}$ to $\boldsymbol{p}\in\mathcal{P}^{n}$. NCMI trains the DNN by encouraging the intra-class concentration and inter-class separation of the output distribution cluster.
@@ III-A Notation
Scalars are denoted by non-bold letters (e.g. $\beta$), vectors by bold lowercase letter (e.g. $\boldsymbol{a}$), the $i$-th entry of a vector $\boldsymbol{a}$ is written as $\boldsymbol{a}_{i}$. We denote $\mathcal{P}^{n}$ as the set of all $n$-dimensional probability distributions. For any two probability distributions $\boldsymbol{p},\boldsymbol{q}\in\mathcal{P}^{n}$, the Kullback–Leibler (KL) divergence is defined as
$D(\boldsymbol{p}\|\boldsymbol{q})=\sum_{i=1}^{n}\boldsymbol{p}_{i}\ln\frac{\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}},$ (1)
where $\ln$ denotes the logarithm with base $e$, write the CE of the one-hot probability distribution corresponding to $y$ and $\boldsymbol{q}$ as $H(y,\boldsymbol{q})=-\ln\boldsymbol{q}_{y}$. Let $\mathcal{S}$ be a collection of probability-simplex-valued functions on $\mathbb{R}^{n}$, i.e.
$\mathcal{S}\subseteq\{\,\sigma:\mathbb{R}^{n}\to\mathcal{P}^{n}\,\},$ (2)
specifically, we define normalized sigmoid function (NSF) $\sigma^{NSF}$ and softmax function $\sigma^{SM}$ as
$\sigma^{NSF}(\boldsymbol{z})_{j}$ $=\frac{\phi(\boldsymbol{z}_{j})}{\sum_{i=1}^{n}\phi(\boldsymbol{z}_{i})},\penalty 10000\ \penalty 10000\ \phi(\boldsymbol{z}_{i})=\frac{1}{1+e^{-\boldsymbol{z}_{i}}};$ (3)
$\sigma^{SM}(\boldsymbol{z})_{i}$ $=\frac{e^{\boldsymbol{z}_{j}}}{\sum_{i=1}^{n}e^{\boldsymbol{z}_{i}}},\text{ where }\boldsymbol{z}\in\mathbb{R}^{n}.$ (4)
Given a multi-class classification dataset $\mathcal{D}$, let $\mathcal{D}^{y}\subseteq\mathcal{D}$ denote the subset of samples labeled $y$.
@@ III-B Modeling Classification as a Markov Chain
In a classification task with $c$ classes, a DNN $f$ and a $\sigma$ could be regarded as a mapping $(\sigma\circ f):\boldsymbol{x}\rightarrow\boldsymbol{p}$, where $\boldsymbol{x}\in\mathbb{R}^{d}$ is an input, and $\boldsymbol{p}\in\mathcal{P}^{n}$ is the output probability distribution. Usually, $n=c$ when we use CE as the surrogate loss. Following [10619241], we can model the classification task as a three-state Markov chain, as depicted in Figure˜1. As shown in [CMI], we empirically quantify the concentration of DNN’s output by
$I(X;\mathcal{P}|Y)$ $=\sum_{y}P_{Y}(y)\sum_{\boldsymbol{x}}P_{X|Y}(\boldsymbol{x}|y)\Big[\sum_{i=1}^{n}\boldsymbol{p}_{i}\ln\frac{\boldsymbol{p}_{i}}{\boldsymbol{s}_{i}^{y}}\Big],$ (5)
$\text{where }\boldsymbol{s}_{i}^{y}$ $\triangleq\frac{1}{|\mathcal{D}^{y}|}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}\boldsymbol{p},\penalty 10000\ \text{for }y\in Y.$ (6)
@@
$\Gamma=\sum_{v\in Y}\sum_{\boldsymbol{x}\in\mathcal{D}^{v}}I_{\{v\neq y\}}P_{X|Y}(\boldsymbol{x}|v)\sum_{i=1}^{n}\boldsymbol{p}_{i}\ln\frac{\boldsymbol{s^{v}_{i}}}{\boldsymbol{p_{i}}}$ (7)
Ideally, we want $I(X;\mathcal{P}|Y=y)$ to be small while keeping $\Gamma$ large. This leads us to consider the ratio between $I(X;\mathcal{P}|Y=y)$ and $\Gamma$.
@@
We refer to $\hat{I}(X;\mathcal{P}|Y)$ as the normalized conditional mutual information (NCMI). To examine how NCMI relates to classification performance, we compute NCMI and top-1 accuracy for pretrained ResNet variants on the ImageNet validation set, as seen in the Figure˜2. We observe a clear inverse linear relationship: models with lower NCMI achieve higher accuracy, with a Pearson correlation coefficient exceeding 0.997-0.997. This suggests that, for a fixed DNN family, improving performance is associated with simultaneously reducing both the error rate and the model’s NCMI during training.
Motivated by this observation, in the next section, we demonstrate that NCMI per se suffices for training DNN classifiers.
@@ IV Methodology
Previous discussions suggest a new surrogate loss for training DNN-based classifiers. Specifically, in the learning process, instead of minimizing the CE surrogate loss, which pulls the output toward predefined distributions, we aim to minimize $\hat{I}(X;\mathcal{P}|Y=y)$. The algorithm for minimizing the novel surrogate loss is outlined below.
@@ IV-A Training DNN by minimizing NCMI
$\min_{\boldsymbol{\theta}}\frac{\sum_{y}P_{Y}(y)\sum_{\boldsymbol{x}}P_{X|Y}(\boldsymbol{x}|y)\Big[\sum_{i=1}^{n}\boldsymbol{p}_{i}\ln\frac{\boldsymbol{p}_{i}}{\boldsymbol{s}_{i}^{y}}\Big]}{\sum_{v\in Y}\sum_{\boldsymbol{x}\in\mathcal{D}^{v}}I_{\{v\neq y\}}P_{X|Y}(\boldsymbol{x}|v)\sum_{j=1}^{n}\boldsymbol{p}_{j}\ln\frac{\boldsymbol{s}^{v}_{j}}{\boldsymbol{p_{j}}}}$ (9)
We notice that the objective in Equation˜9 is not amenable to parallel computation via GPU due to the dependency of $\hat{I}(X;\mathcal{P}|Y)$ on the centroid $\boldsymbol{s^{y}}$ of each cluster corresponding to $Y=y$ (see Equation˜6). To overcome this, we introduce a dummy distribution $\boldsymbol{q^{y}}\in\mathcal{P}^{n}$ for each $y\in[C]$ and convert it into a double minimization problem.
$\min_{\boldsymbol{\theta}}\frac{\sum_{y}P_{Y}(y)\sum_{\boldsymbol{x}}P_{X|Y}(\boldsymbol{x}|y)\Big[\sum_{i=1}^{n}\boldsymbol{p}_{i}\ln\frac{\boldsymbol{p}_{i}}{\boldsymbol{s}_{i}^{y}}\Big]}{\sum_{v\in Y}\sum_{\boldsymbol{x}\in\mathcal{D}^{v}}I_{\{v\neq y\}}P_{X|Y}(\boldsymbol{x}|v)\sum_{i=1}^{n}\boldsymbol{p}_{i}\ln\frac{\boldsymbol{s^{v}_{i}}}{\boldsymbol{p_{i}}}}\equiv$
$\min_{\boldsymbol{q}^{v},v\in[C]}\min_{\boldsymbol{\theta}}\frac{\sum_{y}P_{Y}(y)\sum_{\boldsymbol{x}}P_{X|Y}(\boldsymbol{x}|y)\Big[\sum_{i=1}^{n}\boldsymbol{p}_{i}\ln\frac{\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}^{y}}\Big]}{\sum_{v\in Y}\sum_{\boldsymbol{x}\in\mathcal{D}^{v}}I_{\{v\neq y\}}P_{X|Y}(\boldsymbol{x}|v)\sum_{j=1}^{n}\boldsymbol{p}_{j}\ln\frac{\boldsymbol{q^{v}_{j}}}{\boldsymbol{p_{j}}}}$ (10)
By reformulating the single minimization problem as a double minimization problem, Equation˜10 suggests an alternating algorithm, in which we use gradient descent to minimize the objective function with respect to the model’s parameters $\boldsymbol{\theta}$ and centroids $\boldsymbol{q}^{v},v\in[C]$ iteratively.
@@
9:  Calculate CMI according to Equation˜5
10:# $\textbf{CMI}.\text{shape: }[B,1]$
[…]
12:# $\boldsymbol{\Gamma}.\text{shape: }[B,C]$
13:  $\text{optimizer}_{\boldsymbol{\xi}}.\text{zero\_grad}(),\penalty 10000\ \text{optimizer}_{\boldsymbol{\theta}}.\text{zero\_grad}()$
14:  $\text{loss}\leftarrow(\textbf{CMI}/\boldsymbol{\Gamma}).\text{mean()}$
15:  $\text{loss}.\text{backward()}$
16:  $\text{optimizer}_{\boldsymbol{\xi}}.\text{step}(),\penalty 10000\ \text{optimizer}_{\boldsymbol{\theta}}.\text{step}()$
17:end for

### v3

@@ I Introduction
Deep neural classifiers are typically trained by mapping an input $\boldsymbol{x}$ to a feature vector $\boldsymbol{z}=f_{\theta}(\boldsymbol{x})\in\mathbb{R}^{m}$, followed by a linear classifier $h$ whose softmax produces a distribution on the class simplex. In this de facto setting, cross-entropy (CE) is used as a principled surrogate loss which optimizes class probabilities in a $c$-dimensional probability space determined by the logits.
@@
A large body of work augments CE with additional regularizers [ranasinghe2021orthogonal, hui2023cut, CMI] or proposes ad-hoc alternatives [hui2021evaluation, khosla2020supervised] that empirically compete with CE. However, these approaches typically remain CE-centric; their performance depends on CE, while the auxiliary terms are ineffective on their own, require extensive hyperparameter tuning, or rely on very large batch sizes to realize their gains.
Furthermore, much of the CE literature implicitly operates in an overparameterized regime in which the penultimate feature dimension $m$ is comparable to (or larger than) the number of classes $c$. When $m$ is constrained—e.g., due to compact embeddings for memory/latency or architectures that intentionally keep $m$ small—the terminal linear map induces a low-rank family of class-score vectors whose rank is at most $m$, which becomes a geometric limitation when $c\gg m$ [NEURIPS2019_78f7d96e]. Closely related expressiveness limits of dot-product–Softmax parameterizations have been formalized as the Softmax bottleneck: if the intrinsic rank of the target log-probability matrix exceeds the embedding dimension, no choice of parameters can represent the true conditional distributions [yang2018breaking]. In such constrained-dimension regimes, the usual “CE-on-class-probabilities” viewpoint provides limited guidance on how to shape the representation itself, and the separation intuitions that hold in high-dimensional settings no longer directly apply.
In this paper, we apply a feature-induced simplex formulation. We view a DNN together with a normalized sigmoid function (NSF) as $(\sigma^{NSF}\circ f_{\theta}):\boldsymbol{x}\to\boldsymbol{p}\in\mathcal{P}^{m}$, and inference can be performed via centroid comparison in KL divergence. Our goal is to learn simplex outputs that are (i) highly concentrated within each class and (ii) well separated across classes. Following the information-theoretic modeling in [CMI], we represent classification as a Markov chain and quantify the concentration of the output distribution using the conditional mutual information (CMI) between the input and output, conditioned on the ground-truth label, and we quantify separation with $\Gamma$ (see Section˜III-B) and define the normalized conditional mutual information (NCMI) as the ratio between the CMI and $\Gamma$. Compared with [CMI], where NCMI enters the learning problem as a constraint, we train the DNN classifier directly by minimizing NCMI.
Beyond conceptual alignment, NCMI is practically effective: across pretrained ResNet variants, NCMI exhibits a near-linear inverse relationship with top-1 accuracy, suggesting it is an informative performance surrogate. We further develop an efficient alternating minimization procedure for NCMI optimization. Finally, because mapping features to a simplex can suffer from single-mode collapse (e.g., softmax-induced degeneracy), we incorporate feature centering and a NSF to stabilize training and prevent collapse.
Contributions: (i) an information-theoretic surrogate loss (NCMI) tailored to DNN-based classifiers; (ii) an efficient alternating algorithm for minimizing NCMI; and (iii) comprehensive validation on natural-image and whole-slide imaging benchmarks.
@@ II Related work
Another line of work improves classification accuracy by modifying CE. Focal Loss [lin2017focal] down-weights well-classified samples via a power transformation so training emphasizes hard instances. PolyLoss [leng2022polyloss] reframes standard classification losses as polynomial expansions. Hui et al. propose SquareLoss [hui2021evaluation], and empirically found that squared loss performs on par with or even outperforms CE on modern DNNs. Supervised contrastive learning (SupCon) [khosla2020supervised] pulls together same-class embeddings and pushes apart different-class embeddings, followed by a linear classifier trained on the frozen features. Further, to mitigate overconfident predictions, label smoothing (LS) [szegedy2016rethinking] softens the one-hot targets, which can inadvertently produce compact class clusters [muller2019does]. AntiClass [kat2025anticlasses] replaces the one-hot target with a one-cold target to mitigate neural collapse.
In contrast, this paper studies the surrogate loss for classification through the lens of information geometry. We view a DNN as a mapping from $\boldsymbol{x}\in\mathbb{R}^{d}$ to $\boldsymbol{p}\in\mathcal{P}^{n}$. We trains the DNN by encouraging the intra-class concentration and inter-class separation of the output distribution cluster by directly minimizing NCMI.
@@ III-A Notation
Scalars are denoted by non-bold letters (e.g. $\beta$), vectors by bold lowercase letter (e.g. $\boldsymbol{a}$), the $i$-th entry of a vector $\boldsymbol{a}$ is written as $\boldsymbol{a}[i]$. We denote $\mathcal{P}^{m}$ as the set of all $m$-dimensional probability distributions. For any two probability distributions $\boldsymbol{p},\boldsymbol{q}\in\mathcal{P}^{m}$, the Kullback–Leibler (KL) divergence is defined as
$D(\boldsymbol{p}\|\boldsymbol{q})=\sum_{i=1}^{m}\boldsymbol{p}[i]\ln\frac{\boldsymbol{p}[i]}{\boldsymbol{q}[i]},$ (1)
where $\ln$ denotes the logarithm with base $e$, the CE is defined as
$H(\boldsymbol{p},\boldsymbol{q})=\sum_{i=1}^{m}\boldsymbol{p}[i]\ln\boldsymbol{q}[i],$ (2)
write the CE of the one-hot probability distribution corresponding to $y$ and $\boldsymbol{q}$ as $H(y,\boldsymbol{q})=-\ln\boldsymbol{q}_{y}$. We denote $\mathds{1}$ as the indicator function, and define normalized sigmoid function (NSF) $\sigma^{NSF}$ and softmax function $\sigma^{SM}$ as
$\sigma^{NSF}(\boldsymbol{z})[j]$ $=\frac{\phi(\boldsymbol{z}[j]}{\sum_{i=1}^{m}\phi(\boldsymbol{z}[i])},\penalty 10000\ \penalty 10000\ \phi(\boldsymbol{z}[i])=\frac{1}{1+e^{-\boldsymbol{z}[i]}};$ (3)
$\sigma^{SM}(\boldsymbol{z})[j]$ $=\frac{e^{\boldsymbol{z}[j]}}{\sum_{i=1}^{m}e^{\boldsymbol{z}[i]}},\text{ where }\boldsymbol{z}\in\mathbb{R}^{m}.$ (4)
Given a multi-class classification dataset $\mathcal{D}$, let $\mathcal{D}^{y}\subseteq\mathcal{D}$ denote the subset of samples labeled $y$, and $\boldsymbol{c_{x}}$ denote the label of sample $\boldsymbol{x}$.
@@ III-B Modeling Classification as a Markov Chain
In a classification task with $c$ classes, a DNN $f$, a linear classifier $h$ and a $\sigma$ could be regarded as a mapping $(\sigma\circ h\circ f):\boldsymbol{x}\to\boldsymbol{p_{x}}$, where $\boldsymbol{x}$ is an input, and $\boldsymbol{p_{x}}\in\mathcal{P}^{n}$ is the output probability distribution. Usually, $n=c$ when we use CE as the surrogate loss. Following [10619241], we can model the classification task as a three-state Markov chain, as depicted in Figure˜1. As shown in [CMI], we empirically quantify the concentration of DNN’s output by
$I(X;\mathcal{P}|Y)$ $=\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}D(\boldsymbol{p_{x}}\|\boldsymbol{s}^{y}),$ (5)
$\text{where }\boldsymbol{s}^{y}$ $\triangleq\frac{1}{|\mathcal{D}^{y}|}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}\boldsymbol{p_{x}},\penalty 10000\ \text{for }y\in Y.$ (6)
@@
$\Gamma=\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}D(\boldsymbol{s^{\boldsymbol{c_{x}}}}\|\boldsymbol{p_{z}})$ (7)
Ideally, we want $I(X;\mathcal{P}|Y)$ to be small while keeping $\Gamma$ large. This leads us to consider the ratio between $I(X;\mathcal{P}|Y)$ and $\Gamma$.
@@
We refer to $\hat{I}(X;\mathcal{P}|Y)$ as the normalized conditional mutual information (NCMI). To study its relationship with classification performance, we discard the fully connected classifier of each pretrained ResNet and apply softmax to the penultimate feature vector to obtain a probability vector, from which we compute NCMI according to Equation˜8. For top-1 evaluation, we form class centroids $s^{y}$ as in Equation˜6 and classify each sample by minimum KL-divergence assignment in the induced probability space. Results on the ImageNet validation set are reported in the Figure˜2. We observe a clear inverse linear relationship: models with lower NCMI achieve higher accuracy, with a Pearson correlation coefficient exceeding 0.997-0.997. This suggests that, for a fixed DNN family, improving performance is associated with simultaneously reducing both the error rate and the model’s NCMI during training. Motivated by this observation, in the next section, we demonstrate that NCMI per se suffices for training DNN classifiers.
@@ IV Methodology
Previous discussions suggest a new surrogate loss for training DNN-based classifiers. Specifically, in the learning process, instead of minimizing the CE surrogate loss, which pulls the output toward predefined distributions, we aim to minimize $\hat{I}(X;\mathcal{P}|Y)$. The algorithm for minimizing the NCMI is outlined below.
@@ IV-A Training DNN by minimizing NCMI
$\min_{\boldsymbol{\theta}}\frac{\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}D(\boldsymbol{p_{x}}\|\boldsymbol{s}^{y})}{\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}D(\boldsymbol{s^{\boldsymbol{c_{x}}}}\|\boldsymbol{p_{z}})}$ (9)
We notice that the objective in Equation˜9 is not amenable to parallel computation via GPU due to the dependency of $\hat{I}(X;\mathcal{P}|Y)$ on the centroid $\boldsymbol{s}^{y}$ of each cluster corresponding to $Y=y$ (see Equation˜6). To overcome this, we introduce a dummy distribution $\boldsymbol{q}^{y}\in\mathcal{P}^{n}$ for each $y\in[C]$ and convert it into a double minimization problem, as shown in the following theorem, which will be proved in the Appendix.
Theorem 1
For any DNN: $\boldsymbol{x}\to\boldsymbol{p}$,
$\min_{\boldsymbol{\theta}}\frac{\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}D(\boldsymbol{p_{x}}\|\boldsymbol{s}^{y})}{\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}D(\boldsymbol{s^{\boldsymbol{c_{x}}}}\|\boldsymbol{p_{z}})}\equiv$
$\min_{\boldsymbol{q}^{v},v\in[C]}\min_{\boldsymbol{\theta}}\bigg[\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}}D(\boldsymbol{p_{x}}\|\boldsymbol{q}^{y})\bigg]/$
$\bigg[\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}\Big[H(\boldsymbol{p_{x}},\boldsymbol{p_{z}})-H(\boldsymbol{p_{x}},\boldsymbol{q}^{\boldsymbol{c_{z}}})\Big]\bigg]$ (10)
By reformulating the single minimization as a double minimization, Equation˜10 suggests an alternating algorithm. With the centroids $\boldsymbol{q}^{v},v\in[C]$ fixed, we update the DNN parameters $\theta$ by stochastic gradient descent (SGD) on the mini-batch objective. Then with $\theta$ fixed, the exact minimizer for each $\boldsymbol{q}^{v}$ is the class centroid of the output probability vectors associated with label $v$. Computing these centroids in closed form after every $\theta$-update require dataset-wide aggregation. To keep training efficient, we instead update centroids by SGD. Concretely, we parameterize the centroid $\boldsymbol{q}^{v},v\in[C]$ by an unconstrained vector $\boldsymbol{\xi}^{v}$ and map it to the simplex using softmax, $\boldsymbol{q}^{v}=\sigma^{NSF}(\boldsymbol{\xi}^{v})$; with $\theta$ fixed, we perform one SGD step on $\boldsymbol{\xi}^{v},v\in[C]$ using the same mini-batch loss.
@@
9:  Calculate NCMI according to Equation˜10
10:  $\text{optimizer}_{\boldsymbol{\xi}}.\text{zero\_grad}(),\penalty 10000\ \text{optimizer}_{\boldsymbol{\theta}}.\text{zero\_grad}()$
11:  $\text{{NCMI}}.\text{backward()}$
12:  $\text{optimizer}_{\boldsymbol{\xi}}.\text{step}(),\penalty 10000\ \text{optimizer}_{\boldsymbol{\theta}}.\text{step}()$
13:end for